OpenStudy (anonymous):

7x-17/8=4 solve for x

OpenStudy (anonymous):

x=109/7

OpenStudy (anonymous):

7x-17/8=4 u add 17/8 on both sides 7x-17/8+17/8=4+17/8 U HAVE 2 GET A COMMON DENOMINATOR SO THAT U COULD SUBTRACT A FRACTION THEREFORE, 7X=32/8+17/8 7X=49/8 THEN U KNOW THAT U THEN DIVIDE BY 7 ON BOTH SIDES SO THEN SINCE U HAVE 49/8/7 ULL HAVE TO MULTIPLY BY THE RECIPRICOL OF SEVEN SO ULL REALLY X 49/8 AND 1/7 THEN U COULD REDUCE ACROSS THE 49 AND SEVEN SO IN THE END ULL END UP W/ 7/8 I HOPE THAT HELPED! :)

OpenStudy (anonymous):

why yes yes it did

OpenStudy (anonymous):

LOL K. UR WELCOME